OpenStudy (anonymous):

7x+4y<28 How would I solve this?

OpenStudy (danjs):

solve for what , y ?

OpenStudy (anonymous):

yes, sorry I forgot to include that

OpenStudy (firekat97):

7x + 4y < 28 Leave y on the LHS so you can make the equation in respect to y and simplify. 4y < 28 - 7x Divide both sides by 4, to get y alone y = (28 - 7x)/4 simplify y = 7 - (7/4)x